OpenStudy (rayford):

solve for x 3/x-2/3=5/6 x = 1 x = 2 x = 3 x = 6

OpenStudy (anonymous):

how do you think it will happen?

OpenStudy (rayford):

well..that's why i asked. i really needed help

OpenStudy (anonymous):

i meant have you solved anything by yourself?

OpenStudy (rayford):

i tried but it was just failure..

OpenStudy (anonymous):

show me what you did

OpenStudy (rayford):

i found the lcm then multiplied both sides and ended up with x=3...or answer choice [c]

OpenStudy (rayford):

is this right??

OpenStudy (anonymous):

3/x-2/3=5/6 now shift 2/3 on the other side 3/x=5/6+2/3 now take the LCM on the other side which will be 6 3/x=5+4/6 3/x=9/6 now cross multiply 6*3=9*x 9x=18 x=18/9 x=2

OpenStudy (anonymous):

your procedure was right but you got the wrong answer

OpenStudy (anonymous):

did you get it now?

OpenStudy (rayford):

yes! i see what i was doing wrong!

OpenStudy (anonymous):

great :D i am glad i helped you

OpenStudy (anonymous):

what wrong was you doing btw?

OpenStudy (rayford):

i wasn't shifting to the other side..

OpenStudy (rayford):

however i have 1 more...may you help with this one?

OpenStudy (anonymous):

sure why not :D

OpenStudy (rayford):

Solve for x: 1/2+4/2x=x+4/10

OpenStudy (anonymous):

again shift 1/2 to the other side then take LCM an then cross multiply tell me what you get?

OpenStudy (rayford):

i'm stuck..on the cross multiply part

OpenStudy (rayford):

nevermind i found it! it was -4,-2

OpenStudy (anonymous):

its -4 and 5

OpenStudy (rayford):

i'm sorry it was a typo that's wht i meant ..

OpenStudy (anonymous):

great then :D

OpenStudy (rayford):

thank you for everything..

OpenStudy (anonymous):

no problemo !